Circular No. 8027

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002jl Further to IAUC 8026, W. M. Wood-Vasey et al. report the discovery

of an apparent supernova (mag 19.3) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Nov. 19 UT. Confirming images were made on Nov. 23 (mag 19.2), 26 (mag 19.3), and 30 (mag 19.3). The new object is located at α = 0^{h}20^{m}13^{s}.95, δ = −0^{o}08′20′′.7 (equinox 2000.0), which is within 5′′ of two different neighboring extended sources. No point source is visible at this location in NEAT images taken with the Haleakala 1.2-m telescope on Nov. 6 (limiting mag 20.5 at S/N = 3).

M. Hamuy, Carnegie Observatories, reports that J. Maza (University of Chile) obtained a spectrum (380–930 nm) of SN 2002jl with the Las Campanas Dupont 2.5-m telescope (+ WFCCD) on Dec. 2.13 UT, finding it to be a type-II supernova caught at an early stage. The spectrum is characterized by a blue continuum with Hα, Hβ, and He I 587.6-nm lines exhibiting P-Cyg profiles. An approximate redshift of 0.064 can be derived from weak, narrow emission lines superimposed on the supernova spectrum.

SUPERNOVAE 2002ia, 2002ib, 2002jf, 2002jh Hamuy also reports that spectroscopy (range 320–920 nm), by Maza

and M. Phillips (Carnegie Observatories) with the Baade 6.5-m telescope (+ Boller & Chivens spectrograph) at Las Campanas, shows that SNe 2002ia, 2002ib, 2002jf, and 2002jh are all type-Ia supernovae, with SNe 2002ia (Nov. 29.13 UT) and 2002jf (Nov. 28.19) ∼ 2–3 weeks past maximum, and SNe 2002ib (Nov. 29.10) and 2002jh (Nov. 28.23) ∼ 10 days past maximum. The spectra reveal the following redshifts: SN 2002ia, ≈ 0.072; SN 2002ib, 0.068 (from superimposed narrow emission lines); SN 2002jf, 0.079 (measured from the host galaxy); SN 2002jh, ≈ 0.048.

PKS 0736+017 S. D. Clements, Florida Institute of Technology, writes that CCD ob-

servations during Dec. 5.281–5.549 UT with the SARA 0.9-m telescope on Kitt Peak show this flat-spectrum radio quasar to be in a high optical state, with mean V = 15.16 ± 0.01 and R = 14.61 ± 0.04 (in its quiescent state, V = 16.4 and R = 15.7). In addition, PKS 0736+017 was observed to exhibit microvariability throughout the night, varying by 0.33 mag in V and 0.43 mag in R. Last year, this source exhibited unusual oscillatory behavior while in its high state. Follow-up observations by observers at all wavelengths are encouraged.

2002 December 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green